OpenStudy (anonymous):

check my answers please solve the following equation. 2(h - 8) - h = h - 16 h = 8 h = -8 infinitely many solutions no solution

OpenStudy (anonymous):

answer no solution

OpenStudy (austinl):

\(2(h-8)-h=h-16\) \(2h-16-h=h-16\) \(h-16=h=16\) \(0=0\) Infinitely many solutions!

OpenStudy (anonymous):

oh well th would explain why I couldn't figure it out I think I did it wrong but im almost positively positive that 5 + 7x = 11 + 7x has no solutions am I right?

OpenStudy (austinl):

\(\color{red}{\text{Correction}}\) \(h-16=h-16\)

OpenStudy (austinl):

\(5+7x=11+7x\) \(5-11=7x-7x\) \(-6\ne0\) Correct!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

ok which equation ahs no solution? 8 - (5v + 3) = 5v - 5 has a solution 3m - 6 = 5m + 7 - m has a solution 3w + 4 - w = 5w - 2(w-2) has a solution 7y + 9 = 7y - 6 no solution right

OpenStudy (austinl):

Oh dear, that is a bit of work right there. I would just say solve them all out for the variable. If you have something where "variable"="number" there is a solution. 0=1 no solution 0=0, 1=1, etc. infinitely many solutions.

OpenStudy (anonymous):

so am I right like you only have to do the last one and if it has no solution then its most likely right

OpenStudy (austinl):

The last one, is indeed no solution. \(0\ne-15\)

OpenStudy (anonymous):

whoot cool!

OpenStudy (anonymous):

thx a lot for your help I really appreciate it I wasn't sure about a couple of those :D

OpenStudy (austinl):

No problem, glad to help!